OpenStudy (anonymous):

Tim has 3 cats and 5 dogs. He gives two random animals away to his neighbour. How big is the probability that both of the random animals turn out to be cats.

OpenStudy (dumbcow):

\[\frac{3}{8}*\frac{2}{7} = \frac{3}{28}\]

OpenStudy (dumbcow):

or using combinations \[\frac{3C2}{8C2} = \frac{3}{28}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

What about the probability, that both of the animals are either cats or both of them are dogs?

OpenStudy (dumbcow):

add the probabilities ...p(both cat) +p(both dog)

OpenStudy (anonymous):

3/8*2/7+5/8*4/7?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

Thanks. :)